# Machine-learned trends in mirror configurations in the Large Plasma Device

### v1

@@ VIII.1 Key achievements
To the authors’ knowledge, this is the first time machine learning has been used to infer specific trends and optimize profiles in magnetized plasmas. Three examples of trend inference were shown in this paper: influence of magnetic geometry on plasma width, changes in the axial and radial profiles with changing discharge voltage, and the relationship of gas puff duration with axial gradient scale length. In addition, the axial profile was optimized by minimizing (or maximizing) the axial standard deviation. There is no other way of simultaneously uncovering many trends or finding optima without using an ML model trained on a diverse dataset. Traditionally, such studies would require extensive scans over grids to map the parameter space, but here it was accomplished with a relatively small amount of data.

### v2

@@ VIII.1 Key achievements
To the authors’ knowledge, this work is one of the first instances machine learning has been used to infer specific trends and optimize profiles in magnetized plasmas. Three examples of trend inference were shown in this paper: influence of magnetic geometry on plasma width, changes in the axial and radial profiles with changing discharge voltage, and the relationship of gas puff duration with axial gradient scale length. In addition, the axial profile was optimized by minimizing (or maximizing) the axial standard deviation. There is no other way of simultaneously uncovering many trends or finding optima without using an ML model trained on a diverse dataset. Traditionally, such studies would require extensive scans over grids to map the parameter space, but here it was accomplished with a relatively small amount of data.
@@ VIII.2 Current limitations
Trends such as the dependence of axial gradient on the gas puff duration (fig. 8) or the effect discharge voltage on x-z profiles (fig. 7), although intuitive, remain unverified. Verification of these trends would increase confidence of model predictions when setting LAPD parameters in future experiments.
@@ X Acknowledgements
This work was performed at the Basic Plasma Science Facility, which is a DOE Office of Science, FES collaborative user facility and is funded by DOE (DE-FC02-07ER54918) and the National Science Foundation (NSF-PHY 1036140).